# Instruction-Guided Scene Text Recognition

### v2

@@ 4.5 Comparison with State-of-the-Art
Figure 6: Success and failure cases. All results are obtained with IGTR-AR trained on Union14M-L. The wrong answers are marked in red.
@@ 4.6 Qualitative Analyses
To better understand IGTR, we depict some success and failure cases in Fig. 6. Cases (a)-(i) are all IGTR correctly recognized images, which are rather challenging and many previous methods have failed on them. We illustrate case (i) in detail by asking a few questions and presenting the model’s responses, which are all correct. Note that for current STR models, the model itself could not answer these questions without consulting the recognized text. This example shows the merit of our instruction-guided learning. It well understands character attributes thus leads to correct text reasoning. Nevertheless, there also are a few failure cases. For cases (j)-(m), answers to all the asked questions are logically consistent although misrecognized. For example, case (j) has made exactly the same mistake of recognizing the fifth F as P. Furthermore, cases (k), (l) and (m) correspond to severely blurred, low-contrast, and overexposed text, respectively. They are quite challenging even for humans. However, a majority of characters are still correctly inferred for these examples, demonstrating that our instruction-guided learning has evoked powerful character understanding capability.
@@
In Fig. 7 we give four examples of long text recognition, which is also a typical STR challenge. The results show that IGTR-ER can recognize text exceeding 25 characters. This is because IGTR-ER does not encode the absolute position embedding. Therefore it can overcome this character limit. Nevertheless, the right-bottom case indicates that IGTR-ER may trigger the error of repeated recognition if the same sub-string appears multiple times, which constitutes a practical issue to address.
@@
In addition, as shown in the bottom of Fig. 1, attention maps $Attn=Sum(Attn_{1},Attn_{2},\ldots,Attn_{h})$, where $Attn_{i}$ (Eq. 1) is generated by Question to Image Attention and FFN in Fig. 2, can provide a qualitative explanation for IGTR’s superior performance. When performing attribute prediction, the attention maps show that IGTR can accurately pinpoint and focus on characters emphasized in the instructions. This implies that IGTR effectively correlates the instructions and the image content through the proposed CFMM module. The image patterns are well understood thus enabling smooth execution of the related tasks. As a contrast, previous methods struggle with recognizing rarely seen text image patterns (as shown in the upper of Fig. 1).
@@
Figure 7: The results of three recognition pipelines for long text. Incorrect predictions are marked in red and missed predictions are denoted as underlined.
@@ 5 Conclusion
In this paper, we have presented IGTR, an instruction-guided multi-modal paradigm to build accurate, fast and lightweight STR models. To gain these attractive properties, we have extended existing question-answer instruction to $\left\langle condition,question,answer\right\rangle$ triplet, which not only increases diversity and richness of the questions but also allows efficient parallel learning. We have developed lightweight instruction encoder, cross-modal feature fusion module and multi-task answer head. They together contribute to a series of STR models that present leading accuracy on multiple English and Chinese benchmarks. Meanwhile, the size of these models is uniformly 24.1M and they consume 3.98ms-10.3ms on NVIDIA 1080Ti GPU when inferring a text image. Notably, IGTR enjoys the flexibility of instruction sampling, by adjusting the sampling of character attribute prediction and recognition instructions, various IGTR models are constructed with different balances between accuracy and inference speed. Furthermore, a simple and elegant IGTR model is built to better recognize rarely appearing and morphologically similar characters.
This paper has made a meaningful attempt in constructing small and efficient multi-modal models dedicated to a specific task, i.e., STR. There also are many works ahead. One is long text recognition [70]. Long text is prevalent in real-world applications, as current tools may not well separate multiple words in a line and take them as a whole. Our preliminary attempt has observed the problem of repeated recognition when directly using IGTR-ER. How to tackle this problem is worthy of further study. Another more generic and bigger one is small OCR foundation model, which is capable of accurately and efficiently processing multiple OCR-related tasks like text recognition, document understanding, etc. Similar efforts have emerged in other domains recently [72, 77]. We argue that this research would benefit a wide of applications related to OCR.

### v3

@@ 4.6 Distinction with existing pre-training methods
Figure 6: Success and failure cases. All results are obtained with IGTR-AR trained on Union14M-L training set. Wrong answers are marked in red.
@@ 4.7 Qualitative Analyses
To qualitatively understand IGTR, we depict some success and failure cases in Fig. 6. Cases (a)-(i) are all IGTR correctly recognized images, which are rather challenging and many previous methods have failed on them. We illustrate case (i) in detail by asking some questions and presenting the model’s responses, which are all correct. Note that for current STR models, the model itself could not answer these questions without consulting the recognized text. This example shows the merit of our instruction-guided learning. It well understands character attributes thus leads to correct text reasoning. Nevertheless, there also are a few failure cases. For cases (j)-(m), answers to all the asked questions are logically consistent although misrecognized. For example, case (j) has made exactly the same mistake of recognizing the fifth F as P. Furthermore, cases (k), (l) and (m) correspond to severely blurred, low-contrast, and overexposed text, respectively. They are quite challenging even for humans. However, a majority of characters are still correctly inferred for these examples, demonstrating that our instruction-guided learning has evoked powerful character understanding capability.
@@
To assess the extrapolating-based recognition, in Fig. 7 we give four examples of long text recognition, which is also a typical STR challenge. The results show that IGTR-ER can recognize text exceeding 25 characters. This is because IGTR-ER does not encode the absolute position embedding. Therefore it can overcome this character limit. As can be seen, IGTR-ER correctly recognizes three challenging long text cases that IGTR-PR and IGTR-AP both failed. Nevertheless, the right-bottom case indicates that IGTR-ER may trigger the error of repeated recognition if the same sub-string appears multiple times, which constitutes a practical issue to address.
@@
We also inspect IGTR models from the attention map perspective. As shown in the bottom of Fig. 1, IGTR-AR’s attention maps $Attn=Sum(Attn_{1},Attn_{2},\ldots,Attn_{h})$, where $Attn_{i}$ (Eq. 1) generated by Question to Image Attention and FFN in Fig. 2 is visualized, can provide a qualitative explanation of its superior performance. When performing attribute prediction, the attention maps show that IGTR-AR can accurately pinpoint and focus on characters emphasized in the questions. This implies that IGTR-AR effectively correlates the instructions and the image content through the proposed CFMM module. The image patterns are well understood thus enabling smooth execution of the related tasks. As a contrast, previous methods are localized less satisfactory and struggle with recognizing rarely seen text image patterns (as shown in the upper of Fig. 1).
@@
Figure 7: Results of three recognition pipelines on long text. Incorrect predictions are marked in red and missed predictions are denoted by _. PR, AR, and ER are IGTR-PR, IGTR-AR, and IGTR-ER, respectively.
4.8 More Discussion
For modern text recognizers, recent research focus mostly lies in enhancing recognition performance through various means such as extracting discriminative visual features, incorporating linguistic information, etc. We delve into the distinctions between IGTR and several popular STR models, including SVTR [5], NRTR [10], ABINet [3], and PARSeq [11], from this perspective. Specifically, SVTR is a purely visual model, where the superior recognition performance stems primarily from enhanced visual features. NRTR introduced linguistic clues by leveraging previously decoded characters, which matches the AR-based recognition instruction in Tab. II. PARSeq introduced a learning pattern where a subset of characters is provided, and the model deduces the remaining characters within the text. This learning pattern is encompassed by specific conditions and questions in IGTR, e.g., the first and fourth conditions in Tab. I, the Search status question, etc. ABINet employs a language model-based character double-check mechanism, akin to the Re-identification instruction in Tab. II. On the other hand, IGTR adopts SVTR-B as its visual backbone, and devises a diverse set of instructions to model character-central attributes comprehensively. By doing so, the linguistic information is fully integrated, while the STR models just discussed, from the linguistic modeling point of view, can be viewed as only utilizing distinct subsets of IGTR instructions. This explains why IGTR achieves superior recognition performance. Moreover, the instruction-guided training in IGTR also enjoys great flexibility, targeted strengthened models can be easily constructed by simply adjusting the instructions sampling, which is difficult to achieve by existing STR models. This showcases another advantage of IGTR in terms of adaptability.
@@ 5 Conclusion
In this paper, we have presented IGTR, an instruction-guided multi-modal paradigm to build accurate, fast and lightweight STR models. To gain these attractive properties, we have extended traditional question-answer instruction to $\left\langle condition,question,answer\right\rangle$ triplet, which not only increases the diversity and richness of the instructions but also allows efficient parallel learning. We have developed a lightweight instruction encoder, a cross-modal feature fusion module and a multi-task answer head. They together generate a novel architecture with sufficient cross-modal feature interactions, contributing to IGTR models that present leading accuracy on various English and Chinese benchmarks. Notably, IGTR enjoys the flexibility of instruction sampling, by using different text recognition instructions, IGTR models following different recognition pipelines are constructed and achieve different trade-offs between accuracy and inference speed. Moreover, by adjusting the sampling of attribute prediction instructions, a targeted strengthened IGTR model can be built, which better recognizes rarely appearing and morphologically similar characters. It is worth noting that the size of IGTR models is uniformly 24.1M, and they consume 3.98 ms-10.3 ms on average using one NVIDIA 1080Ti GPU when inferring a text instance. Both are appealing properties in resource-constrained applications.
This paper has made a meaningful attempt in developing small and efficient multi-modal models dedicated to a specific task, i.e., STR. There also are many works ahead. One is long text recognition [73]. Long text is prevalent in real-world applications, as current tools may not correctly separate multiple words in a line and text recognizers should take them as a whole. Our preliminary attempt has observed the problem of repeated recognition when directly using IGTR-ER. How to tackle this problem is worthy of further study. Another critical issue is small OCR foundation model, which is capable of accurately and efficiently processing multiple OCR-related tasks like text recognition, document understanding, etc. This research would benefit a wide array of applications related to OCR. We believe that similar initiatives in other domains [83, 43, 84] can shed light on us.